OpenStudy (anonymous):

whats the pattern?... 2,8,32,128

OpenStudy (anonymous):

Multiply by 4 each time.

OpenStudy (anonymous):

The pattern appears to be: \[ 2^{2n-1}=A_n \]

OpenStudy (anonymous):

or odd powers of 2 2^1, 2^3, 3^5, 2^7

OpenStudy (anonymous):

2^5*

OpenStudy (anonymous):

evaluate. -4+2(-3+7)2

OpenStudy (anonymous):

the odd powers is the same and x4 cause its like 2 x 2^2 , 2^3 x 2^2 etc

OpenStudy (anonymous):

-4+2(-3+7)^2 -4+2(4)^2 -4+2(16) -4+32 28

OpenStudy (anonymous):

thank you! what about this problem? 16/(2+6)* 10

OpenStudy (anonymous):

/ = divide * + multiply

OpenStudy (anonymous):

16/10(2+6) 16/10(8) 16/80 1/5

OpenStudy (anonymous):

/ = division

OpenStudy (anonymous):

cool?

OpenStudy (anonymous):

no its not a fraction..

OpenStudy (anonymous):

nevermind!! sorry (:

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

simplify 3x + 4y -7y + x - 10x

OpenStudy (anonymous):

-3y-9x+3x -3y-6x -3(y+2x)

OpenStudy (anonymous):

simplify -14z^3 + 11z - 8z^2 -9z^3 =7z^2 =5z

OpenStudy (anonymous):

and all the = is suposed to be a +

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

kk 1 sec

OpenStudy (anonymous):

-14z^3-9z^3-8z^2+7z^2+11z+5z -23z^3-z^2+16z z(-23z^2-z+16)

OpenStudy (anonymous):

can you evaluate this equation.. x^2 + 5 - x when x = -6

OpenStudy (anonymous):

x^2+5-x (-6)^2+5-(-6) 36+5+6 47

OpenStudy (anonymous):